# Who Gets Cited Most? Benchmarking Long-Context Language Models on Scientific Articles

### v1

@@ Abstract
This paper introduces SciTrek, a novel question-answering benchmark designed to evaluate the long-context reasoning capabilities of large language models (LLMs) using scientific articles. Current long-context benchmarks often rely on non-scientific texts, focus on simple information retrieval tasks, or employ artificial contexts. SciTrek addresses these limitations by proposing complex questions that require information aggregation and synthesis across multiple full-text scientific articles. Questions and their ground-truth answers are automatically generated by formulating them as SQL queries over a database constructed from article metadata (titles, authors, and references). The SQL operations provide explicit, verifiable reasoning steps for fine-grained error analysis, and the construction process scales to contexts up to 1M tokens with minimal supervision. Extensive experiments on a diverse set of open-weight and proprietary LLMs demonstrate that SciTrek poses a significant challenge as the context length increases, with supervised fine-tuning and reinforcement learning offering only limited gains. Our analysis reveals systematic shortcomings in models’ abilities to perform basic numerical operations and accurately locate specific information in long contexts.111Our dataset can be downloaded from https://github.com/oaimli/SciTrek
1 Introduction
While long-context language models (LCLMs) are highly relevant to scientific tasks, existing benchmarks designed to evaluate their capabilities predominantly focus on non-scientific texts (Maekawa et al., 2025; Bai et al., 2025; Yen et al., 2025; Kuratov et al., 2024). They mostly address simple information retrieval tasks (e.g., Needle-In-A-Haystack; Kamradt 2023; Hsieh et al. 2024) rather than information integration, i.e., reasoning across multiple documents to synthesize information and answer complex questions. Some question-answering benchmarks (Asai et al., 2024; Bai et al., 2025; An et al., 2024) require information integration, but do not provide a typology of the reasoning steps involved, which makes it difficult to determine the specific capabilities being assessed and why a model fails. Another issue relates to scalability and extensibility. When substantial human effort is involved in generating the questions and their answers (e.g., Asai et al. 2024), scaling to longer contexts or different types of questions becomes prohibitive. To better emulate scientific research practice, it is also important to evaluate a range of skills in natural contexts, such as a model’s ability to aggregate or filter information or combinations thereof (Wang et al., 2024c). However, a few benchmarks rely on artificial contexts, often created by extending short-context tasks (e.g., HotpotQA; Yang et al. 2018) with added noise (Kuratov et al., 2024) or by synthesizing the input with LLMs (Maekawa et al., 2025).
In this paper, we create a question-answering benchmark based on scientific articles that aims to alleviate these issues. We assume a scenario where a researcher, possibly during the literature review process, seeks answers to specific questions about the scientific articles they read. As an example, consider the first question in Table 1: “What is the highest number of authors that any single article has?”. Although seemingly simple, it requires aggregation of information across multiple articles. To answer correctly, a hypothetical model would need to count the number of authors for each article within a document collection, and identify the article with the largest number. To generate questions with natural language inputs and explicit reasoning steps, we focus on basic numerical operations such as counting, aggregating, and sorting, applied to key elements of scientific articles: titles, authors, and references. Since these operations correspond to standard database functions, we can automatically generate arbitrarily complex SQL queries over any collection of scientific articles and obtain ground-truth answers without manual annotation by constructing database tables for titles, authors, and references. We take advantage of recent advances in LLM capabilities for SQL understanding and generation (Yu et al., 2018; Li et al., 2023; Hui et al., 2024) to automatically convert SQL queries into corresponding natural language questions.
@@ 1 Introduction
| How many references do articles with exactly two authors have? | SELECT reference_count FROM articles WHERE author_count = 2 | 16 |
@@
| How many articles have been cited by other articles but do not cite any other articles? | SELECT COUNT(*) FROM articles WHERE article_id NOT IN (SELECT article_id_citing FROM citing_cited) AND article_id IN (SELECT article_id_cited FROM citing_cited) |
@@
Our benchmark, SciTrek, contains a test set of 2,121 question-answer pairs on scientific articles and a training set of 19,543 question-answer pairs, with contexts of varying lengths (i.e., 64K, 128K, 512K, and 1M555Throughout the paper, K refers to 1,024 and M to 1,024K.). Table 1 illustrates examples of questions, their corresponding SQL queries, and answers. These questions are relatively superficial, as they do not engage deeply with article content. However, the metadata consists of distillable facts sparsely distributed across article collections, requiring models to retrieve and synthesize information over long contexts. If models struggle with simple filtering and sorting, they are unlikely to succeed on more complex analytical tasks. We use SciTrek to evaluate the capabilities of LCLMs with context windows exceeding 128K tokens. Experimental results reveal that the benchmark poses significant challenges, with models achieving poor performance, even when enhanced with supervised fine-tuning (SFT) and reinforcement learning (RL). We make the following contributions:
@@
We propose a methodology for constructing a long-context question-answering benchmark over natural texts, with explicit reasoning steps approximated via SQL operations that involve information aggregation. The methodology requires minimal human intervention and can be replicated for other domains containing entities that can be structured as database tables.
@@
Extensive experiments reveal that frontier open-weight and proprietary LLMs struggle significantly with this task, especially as input length increases. While SFT and RL improve over baseline models, performance remains limited. All models struggle with questions involving references and citation relationships among the input articles.
@@
By design, SciTrek enables fine-grained analysis of model behavior, offering insight into where and why models fail. Our evaluation of open-weight models reveals systematic shortcomings in counting bibliographic elements and performing basic numerical operations, with error rates increasing as question length grows. Interestingly, models frequently misinterpret compound conditions and struggle with logical constructs involving negation. (see the last question in Table 1).
@@
| Dataset | Natural | Skills | Reason | Scalable | Length | Example Question |
| NeedleBench | ✗ | ✗ | ✗ | ✓ | 128K | What legendary item is hidden on Emerald Island? |
| Ada-LEval | ✗ | ✗ | ✗ | ✓ | 128K | What is the correct order of the segments? |
| BABILong | ✗ | ✗ | ✗ | ✓ | 10M | Where is Mary? |
| HELMET | ✗ | ✓ | ✗ | ✗ | 128K | Who set the fire in one tree hill? |
| LIFBENCH | ✗ | ✓ | ✗ | ✓ | 128K | Retrieve the entry at position 8th in the list. |
| RULER | ✗ | ✓ | ✗ | ✓ | 128K | Find all variables that are assigned the value 12345. |
| OpenScholar | ✓ | ✓ | ✗ | ✗ | 3K∗ | Compile a list of reviews […], and identify the most promising […] |
@@
| LongBench v2 | ✓ | ✓ | ✗ | ✗ | 128K | How long have I been living in my current apartment in Shinjuku? |
| LongMemEval | ✓ | ✓ | ✗ | ✗ | 2M | How many bikes do I currently own? |
| L-Eval | ✓ | ✓ | ✗ | ✗ | 200K | How do I know when I should apply for Medicare? |
@@
| HoloBench | ✗ | ✓ | ✓ | ✓ | 64K | What are the names of wines and their corresponding grape types? |
@@
| MathHay | ✗ | ✓ | ✓ | ✓ | 128K | What is the total number of points scored by LeBron […] combined? |
| Loong | ✓ | ✗ | ✓ | ✓ | 250K | Which company has the highest non-current assets? |
| SciTrek | ✓ | ✓ | ✓ | ✓ | 1M | How many articles have been cited by other articles but do not cite any other articles? |
@@ 2 Related Work
A growing body of work explores the challenges LLMs encounter when processing inputs that exceed their context lengths. Rotary Position Embeddings (RoPE; Su et al. 2024) have emerged as a widely used approach, replacing absolute positional embeddings with rotational transformations of token embeddings. It improves generalization to unseen sequence lengths. Building on RoPE, several extensions introduce rescaling to further enhance performance on substantially longer inputs without the need for retraining (Peng et al., 2024).
Other architectural advances focus on reducing the quadratic cost of the Transformer attention mechanism (Sun et al., 2025). Sparse attention methods, such as LongFormer (Beltagy et al., 2020; Jiang et al., 2024; Fu et al., 2024), compute only selected portions of the full attention matrix, thereby reducing computational cost and accelerating both prefilling and inference. Many contemporary models interleave sparse and full attention layers to balance efficiency and performance (Llama Team, 2025; Yang et al., 2025; Kamath et al., 2025; Dubey et al., 2024). Another popular strategy of reducing the computation required in long-context settings is to limit the number of active model parameters. Group Query Attention (GQA; Ainslie et al. 2023) enables multiple attention heads to share key–value projections, without significant performance loss. The open-weight LLMs used in our experiments adopt combinations of these techniques, complemented by pretraining on long-context tasks (Dubey et al., 2024; Kamath et al., 2025; Yang et al., 2025; Llama Team, 2025).
@@
With growing interest in LCLMs, there has been a surge in benchmarks designed to evaluate their performance. As far as text-based efforts are concerned, initial benchmarks primarily assessed whether models are able to retrieve relevant information from their context, typically through Needle-In-A-Haystack (NIAH) tasks, such as NeedleBench (Li et al., 2024). Building on these, subsequent benchmarks introduced either distractor information to short-context tasks or dispersed relevant information across synthetic contexts, such as Ada-LEval (Wang et al., 2024a), BABILong (Kuratov et al., 2024), HELMET (Yen et al., 2025), LIFBENCH (Wu et al., 2025b), LongBench (Bai et al., 2024), and RULER (Hsieh et al., 2024).
In addition to retrieval, some benchmarks emphasize complex reasoning over long contexts. For example, OpenScholar (Asai et al., 2024) and LongBench v2 (Bai et al., 2025) contains questions that require extensive use of the multiple documents provided as context, while DocFinQA (Reddy et al., 2024) and MedOdyssey (Fan et al., 2025) focus on financial and medical reasoning, respectively. Although not explicitly technical, LongMemEval (Wu et al., 2025a) and L-Eval (An et al., 2024) also requires reasoning over natural, long-context information. However, evaluating model answers against these benchmarks remains challenging, or the reasoning steps required to arrive at correct responses are often opaque, making it difficult to assess failure modes of LCLMs. In contrast, HoloBench (Maekawa et al., 2025) and MathHay (Wang et al., 2024b) better delineate between the types of reasoning required for each question, but rely on unnatural contexts. Loong (Wang et al., 2024c) combines structured reasoning and natural contexts, but tests a very limited number of skills. CURIE (Cui et al., 2025), OpenScholar (Asai et al., 2024) and LongBench v2 (Bai et al., 2025) evaluate reasoning over scientific articles, but their reliance on expert-curated questions and answers poses challenges for scalability to longer contexts or larger datasets.
In designing SciTrek, we aim to retain many of the useful qualities of previous benchmarks (e.g., structured reasoning, scalability) while still proposing a fairly natural and easy task for humans to perform. Although we focus on scientific articles, our methodology extends to other domains with explicit entities and relations. Thanks to the SQL backbone, we are able to construct questions testing various model skills (e.g., aggregation, filtering), and how these manifest themselves through explicit reasoning steps. We compare SciTrek with representative existing benchmarks in Table 2.
@@ 3 The SciTrek Benchmark
In this section, we explain how SciTrek was curated. As mentioned earlier, our benchmark consists of question-answer pairs over scientific articles. Figure 1 illustrates our process: we first gather scientific article collections as contexts corresponding to different lengths (i.e., 64K, 128K, 512K, and 1M); we then create databases representing the article metadata and SQL queries with answers from database execution; finally, we convert the SQL queries into natural language questions.
@@ 3.1 Gathering Scientific Article Collections
SciTrek is constructed using scientific articles from Semantic Scholar666https://www.semanticscholar.org/. To cover diverse topics of articles, we obtain an initial set of seed articles from eight subjects: Computer Science (CS), Economics, Electronic Engineering (EE), Math, Physics, Biology, Finance, and Statistics.777Following the typology of subjects from https://arxiv.org/. For each subject, we select two seed articles with more than 100 citations since 2020. For each article, we retrieve related articles from Semantic Scholar.888Semantic Scholar provides APIs to retrieve scientific articles with metadata including titles, authors, a reference list with articles that the current article cites, and a citation list with articles that cite the current article. Throughout this paper, a reference is a bibliographic entry listed in the reference section of a scientific article, whereas a citation is an in-text acknowledgement of another’s work within the article. To ensure broad coverage, we include two-hop related articles identified via Semantic Scholar’s citation graph based on the reference and citation list. For each seed article, we randomly sample ten first-hop related articles, and for each of these, we further sample five second-hop related articles to form an article cluster. Since full texts are required, we filter out articles without PDFs. This process yields 16 article clusters comprising 662 scientific articles with PDFs across eight subjects (see step I in Figure 1). Finally, we convert the collected PDFs into markdown texts using Marker.999https://github.com/datalab-to/marker
@@
From these clusters, we construct article collections of varying context lengths for our question-answering task. By concatenating the markdown texts, we generate collections with lengths of 64K, 128K, 512K, and 1M tokens.101010We can easily construct collections with arbitrary lengths exceeding 1M tokens. Each collection is initialized with a randomly selected article, and additional articles are added until it reaches a specified length level. We expand the collections using two strategies: (1) random sampling from the clusters, and (2) traversing the citation graph within each cluster (using both depth-first and breadth-first search) to construct collections that preserve citation relations among articles. Each collection contains at least four articles, and no two collections share more than half of their articles. From the 662 scientific articles across 16 clusters, we construct 2,612 article collections spanning all four length levels, including 2,027 generated through random sampling and 585 through graph traversal.

### v2

# Who Gets Cited Most? Benchmarking Long-Context Reasoning on Scientific Articles

@@ Abstract
We introduce SciTrek, a novel question-answering benchmark designed to evaluate long-context reasoning capabilities of large language models (LLMs) using scientific articles. Current long-context benchmarks often focus on simple information retrieval tasks, or employ artificial contexts. SciTrek addresses these limitations by creating benchmark questions that require information aggregation and synthesis across multiple full-text scientific articles. The questions and their ground-truth answers are automatically generated by formulating them as SQL queries over a database constructed from article metadata (i.e., titles, authors, and references). These SQL queries provide explicit, verifiable reasoning processes that enable fine-grained error analysis on model answers, and the data construction scales to contexts of up to 1M tokens with minimal supervision. Experiments on open-weight and proprietary LLMs show that SciTrek poses significant challenges as the context length increases, with supervised fine-tuning and reinforcement learning offering only limited gains. Our analysis reveals systematic shortcomings of frontier LLMs’ ability to effectively perform numerical operations and accurately locate information in long contexts.111Our dataset and code are available from https://github.com/oaimli/SciTrek
@@
| How many references do articles with exactly two authors have? | SELECT reference_count FROM articles WHERE author_count = 2 |
16
@@
| How many articles are cited by other articles but do not cite any other articles? | SELECT COUNT(*) FROM articles WHERE article_id NOT IN (SELECT article_id_citing FROM citing_cited) AND article_id IN (SELECT article_id_cited FROM citing_cited) |
@@
1 Introduction
While long-context language models (LCLMs) are highly relevant to scientific tasks, existing benchmarks for evaluating their capabilities predominantly focus on non-scientific texts (Kuratov et al., 2024; Maekawa et al., 2025; Bai et al., 2025; Yen et al., 2025). They mostly address simple information retrieval tasks (e.g., Needle-In-A-Haystack; Kamradt 2023; Hsieh et al. 2024) rather than information integration, i.e., reasoning across multiple documents to synthesize information and answer complex questions. Some question-answering benchmarks (Asai et al., 2024; An et al., 2024; Bai et al., 2025) require information integration but do not provide a typology of the reasoning steps involved, making it difficult to determine which capabilities are being assessed and why a model fails.
Another issue relates to scalability and extensibility. When substantial human effort is required to generate benchmark questions and their answers (e.g., Asai et al. 2024), scaling to longer contexts or different types of questions becomes prohibitive. To better emulate scientific research practice, it is also important to evaluate a range of skills in natural contexts, such as a model’s ability to aggregate or filter information, and combinations thereof (Wang et al., 2024c). In contrast, a few benchmarks rely on artificial contexts, created by adding noise (Kuratov et al., 2024) to short contexts (e.g., HotpotQA; Yang et al. 2018) or by synthesizing the input with LLMs (Maekawa et al., 2025).
In this paper, we create a question-answering benchmark based on scientific articles that aims to alleviate these issues. We assume a scenario where a researcher, possibly during the literature review process, seeks answers to specific questions when the related scientific articles are provided. As an example, consider the first question in Table 1: “What is the highest number of authors that any single article has?”. Although seemingly simple, it requires aggregation of information across multiple articles. To answer correctly, a hypothetical model would need to count the authors of each article in the collection and select the one with the largest number. We generate questions requiring explicit reasoning by focusing on basic numerical operations such as counting, aggregating, and sorting, applied to core elements of scientific articles: titles, authors, and references. Since these operations correspond to standard database functions, we can automatically generate arbitrarily complex SQL queries over any collection of scientific articles and obtain ground-truth answers without manual annotation by constructing database tables for titles, authors, and references. We exploit recent advances in LLM capabilities for SQL understanding and generation (Yu et al., 2018; Li et al., 2023; Hui et al., 2024) to automatically convert SQL queries into natural-language questions.
@@ 1 Introduction
Our benchmark, SciTrek, contains a test set of 2,121 question-answer pairs on scientific articles and a training set of 19,543 question-answer pairs, with contexts of varying lengths (i.e., 64K, 128K, 512K, and 1M).555In the paper, K refers to 1,024 and M to 1,024K. Table 1 shows examples of questions, their corresponding SQL queries, and answers. While these questions are relatively superficial and do not engage deeply with the domain-specific content, they require models to retrieve and synthesize facts that are sparsely distributed across article collections and span long contexts. If models struggle with simple comparison, filtering, sorting, and aggregation operations, they might be impossible to succeed on more complex analytical tasks (Wolfson et al., 2020, 2025). We use SciTrek to evaluate the reasoning capabilities of LCLMs with context windows exceeding 128K tokens. Experimental results reveal that SciTrek poses significant challenges for these LCLMs. In general, models achieve poor performance, even when enhanced with supervised fine-tuning (SFT) and reinforcement learning (RL). This paper makes the following contributions:
@@
We propose a methodology for constructing a long-context question-answering benchmark over natural texts, where explicit reasoning processes are approximated via SQL operations involving information aggregation. The methodology requires minimal human intervention and generalizes to any domain with entities that can be structured as database tables.
@@
Extensive experiments reveal that frontier open-weight and proprietary LLMs struggle with this task, especially as input length increases. While SFT and RL improve over baseline models, performance remains limited. All models perform poorly on questions involving references and citation relationships among articles.
@@
By design, SciTrek enables fine-grained analysis of model behavior, offering insight into where and why models fail. Evaluation of open-weight models reveals systematic shortcomings in counting bibliographic elements and performing basic numeric operations, with error rates increasing as question length grows; models also frequently misinterpret compound conditions and struggle with logical constructs involving negation (see last question in Table 1).
@@
| Dataset | Natural Context | #SciQs | Reasoning | Scalable | Len | Example Question |
| NeedleBench | ✗ | — | ✗ | ✓ | 128K | What legendary item is hidden on Emerald Island? |
| Ada-LEval | ✗ | — | ✗ | ✓ | 128K | What is the correct order of the segments? |
| BABILong | ✗ | — | ✗ | ✓ | 10M | Where is Mary? |
| HELMET | ✗ | — | ✗ | ✗ | 128K | Who set the fire in one tree hill? |
| LIFBENCH | ✗ | — | ✗ | ✓ | 128K | Retrieve the entry at position 8th in the list. |
| RULER | ✗ | — | ✗ | ✓ | 128K | Find all variables that are assigned the value 12345. |
| OpenScholar | ✓ | 208 | ✗ | ✗ | 3K∗ | Compile a list of reviews […], and identify the most promising […] |
@@
| LongBench v2 | ✓ | 50 | ✗ | ✗ | 128K | How long have I been living in my current apartment in Shinjuku? |
| LongMemEval | ✓ | — | ✗ | ✗ | 2M | How many bikes do I currently own? |
| L-Eval | ✓ | — | ✗ | ✗ | 200K | How do I know when I should apply for Medicare? |
@@
| HoloBench | ✗ | — | ✓ | ✓ | 64K | What are the names of wines and their corresponding grape types? |
@@
| MathHay | ✗ | — | ✓ | ✓ | 128K | What is the total number of points scored by LeBron […] combined? |
| Loong | ✓ | 53 | ✓ | ✓ | 250K | Which company has the highest non-current assets? |
| SciTrek | ✓ | 1,716 | ✓ | ✓ | 1M | How many articles are cited by other articles but do not cite any other articles? |
@@ 2 Related Work
A growing body of work explores the challenges LLMs encounter when processing inputs that exceed their context lengths. Rotary Position Embeddings (RoPE; Su et al. 2024) have emerged as a widely used approach that replaces absolute positional embeddings with rotational transformations of token embeddings, improving generalization to unseen sequence lengths. Building on RoPE, several extensions introduce rescaling to further enhance performance on substantially longer inputs without the need for retraining (Peng et al., 2024).
Other architectural advances focus on reducing the quadratic cost of the Transformer attention mechanism (Sun et al., 2025). Sparse attention methods, such as LongFormer (Beltagy et al., 2020), compute only selected portions of the full attention matrix, thereby reducing computational cost and accelerating both prefilling and inference (Jiang et al., 2024; Fu et al., 2024). Many contemporary models interleave sparse and full attention layers to balance efficiency and performance (Dubey et al., 2024; Llama Team, 2025; Yang et al., 2025; Kamath et al., 2025). Another popular strategy for reducing computation in long-context settings is to limit the number of active model parameters. Group Query Attention (GQA; Ainslie et al. 2023) enables multiple attention heads to share key–value projections, without significant performance loss. The open-weight LLMs used in our experiments combine these techniques, complemented by pretraining on long-context tasks (Dubey et al., 2024; Kamath et al., 2025; Yang et al., 2025; Llama Team, 2025).
@@
The growing interest in LCLMs has driven the development of numerous benchmarks designed to evaluate their performance. Initial benchmarks primarily assessed whether models can retrieve information from their context, typically through Needle-In-A-Haystack (NIAH) tasks, such as NeedleBench (Li et al., 2024). Building on these approaches, subsequent benchmarks either added distractor information to short-context tasks or scattered relevant information across synthetic contexts, such as Ada-LEval (Wang et al., 2024a), BABILong (Kuratov et al., 2024), HELMET (Yen et al., 2025), LIFBENCH (Wu et al., 2025b), LongBench (Bai et al., 2024), and RULER (Hsieh et al., 2024).
In addition to retrieval, some benchmarks emphasize complex reasoning over long contexts. For example, OpenScholar (Asai et al., 2024) and LongBench v2 (Bai et al., 2025) contain questions that require extensive use of multiple documents provided as context, while DocFinQA (Reddy et al., 2024) and MedOdyssey (Fan et al., 2025) focus on financial and medical reasoning, respectively. Although not explicitly technical, LongMemEval (Wu et al., 2025a) and L-Eval (An et al., 2024) also require reasoning over natural, long-context information. However, evaluating model answers against these benchmarks remains challenging, as the reasoning steps required to arrive at correct responses are often opaque, making it difficult to assess failure modes. In contrast, HoloBench (Maekawa et al., 2025) and MathHay (Wang et al., 2024b) better delineate between the types of reasoning required for each question, but rely on unnatural contexts. CURIE (Cui et al., 2025), OpenScholar, and LongBench v2 evaluate reasoning over scientific articles, but their reliance on expert-curated questions and answers poses challenges for scalability to longer contexts or larger datasets.
In designing SciTrek, we aim to retain many of the useful qualities of previous benchmarks (e.g., structured reasoning, scalability) while still proposing a fairly natural and easy task for humans to perform. We compare SciTrek with existing benchmarks in Table 2. Unlike Loong (Wang et al., 2024c), which tests broad long-context comprehension, SciTrek uniquely targets scientific synthesis and structured reasoning at extreme context lengths. Although we focus on scientific articles, our methodology could extend to other domains with explicit entities and relations. Thanks to the SQL backbone, we are able to construct questions testing various model skills (e.g., aggregation, filtering), and how these manifest themselves through explicit reasoning processes.
@@ 3 The SciTrek Benchmark
Figure 1 illustrates how SciTrek was curated: we first gather scientific article collections as contexts corresponding to different lengths (i.e., 64K, 128K, 512K, and 1M); we then create databases representing the article metadata and SQL queries with answers from database execution; finally, we convert the SQL queries into natural language questions. We explain these stages in detail below.
@@ 3.1 Gathering Scientific Article Collections
SciTrek is constructed using scientific articles from Semantic Scholar.666https://www.semanticscholar.org/ To cover diverse topics of articles, we obtain an initial set of seed articles from eight subjects: Computer Science (CS), Economics, Electronic Engineering (EE), Math, Physics, Biology, Finance, and Statistics.777Following the typology of subjects from https://arxiv.org/. For each subject, we select two seed articles with more than 100 citations since 2020. For each article, we retrieve related articles from Semantic Scholar.888Semantic Scholar provides APIs to retrieve scientific articles with metadata including titles, authors, a reference list with articles that the current article cites, and a citation list with articles that cite the current article. Throughout this paper, a reference is a bibliographic entry listed in the reference section of a scientific article, whereas a citation is an in-text acknowledgement of another’s work within the article. To ensure broad coverage, we include two-hop related articles identified via Semantic Scholar’s citation graph based on the reference and citation list. For each seed article, we randomly sample ten first-hop related articles, and for each of these, we further sample five second-hop related articles to form an article cluster. Since full texts are required, we filter out articles without PDFs. This process yields 16 article clusters comprising 662 scientific articles with PDFs across eight subjects (see step I in Figure 1). Finally, we convert the collected PDFs into markdown texts using Marker.999https://github.com/datalab-to/marker
@@
From these clusters, we create article collections of varying context lengths for our question-answering task. By concatenating the markdown texts, we generate collections with lengths of 64K, 128K, 512K, and 1M tokens.101010We can easily construct collections with arbitrary lengths exceeding 1M tokens. Each collection is initialized with a randomly selected article, and additional articles are added until it reaches a specified length level. We expand the collections using two strategies: (1) random sampling from the clusters, and (2) traversing the citation graph within each cluster (using both depth-first and breadth-first search) to construct collections that preserve citation relations among articles. Each collection contains at least four articles, and no two collections share more than half of their articles. From the 662 scientific articles across 16 clusters, we construct 2,612 article collections spanning all four length levels, including 2,027 generated through random sampling and 585 through graph traversal.
3.2 Creating Databases and SQL Queries
Once the article collections are assembled, we construct a database for each (see step II in Figure 1), focusing on a core subset of elements shared across all articles: titles, authors, and references. Based on these key elements, each database comprises three tables: articles, article-author, and citing-cited. The articles table contains metadata such as the title, reference count, and title word count. The article-author table captures author information, including names and their positions in the corresponding author list. The citing-cited table records citation relations among articles. A detailed description is provided in Appendix A. This information is obtained from Semantic Scholar or derived via simple preprocessing, such as splitting titles and counting words. Manual corrections were applied as needed to align with the full-text Markdown articles.